Schedule of Cash Collections of Accounts Receivable Pet Place Supplies Inc., a pet wholesale supplier, was organized on May 1Collected in July Total cash collected
Answer #1
Pet place Supplies Inc
Schedule of cash collection from sales
For the three months Ending July 31
May June July
Receipts from cash sales
May sales on account
Collected in May $310,000*51% = $158,100
Collected in June $310,000*45% = $139,500
Collected in July $310,000*4% = $12,400
June sales on account
Collected in June $420,000*51% = $214,200
Collected in July $420,000*45% = $189,000
July sales on account
Collected in July $560,000*51% = $285,600
Total cash receipts $                                   158,100 $                                   353,700 $                                   487,000

You can reach me over comment box if you have any doubts. Please rate this answer